OpenStudy (anonymous):

Express the complex number in trigonometric form. -3 + 3 square root of 3i.

OpenStudy (loser66):

\(-3+3\sqrt{3}i\) right?

OpenStudy (anonymous):

Yes.

OpenStudy (loser66):

Let wait for Zepdrix, he is better than me hihihi

zepdrix (zepdrix):

lol XD

OpenStudy (loser66):

@zepdrix ,please

OpenStudy (anonymous):

Hehe, its okay. I'm the one coming to you guys afterall.

zepdrix (zepdrix):

So you can write it in the form:\[\Large\rm r(\cos \theta+\mathcal i \sin \theta)\]^That is our trig form. We need to find \(\Large\rm r\) and \(\Large\rm \theta\)

zepdrix (zepdrix):

When we have:\[\Large\rm a+b\mathcal i\]Our radius is given by:\[\Large\rm r=\sqrt{a^2+b^2}\]

zepdrix (zepdrix):

So what do we get for a radius in this case? :D Crunch some numbers!!

OpenStudy (anonymous):

\[-3^{2} + 3\sqrt3^{2} = 0\]

OpenStudy (anonymous):

So, the radius would be 0.

zepdrix (zepdrix):

You got zero for a radius? XD

zepdrix (zepdrix):

\[\Large\rm -3+3\sqrt{3}\mathcal i\] \[\Large\rm r=\sqrt{(-3)^2+(3\sqrt3)^2}\]Don't forget to square the negative sign!!

OpenStudy (anonymous):

Ah right, I forgot my calculator always gives me negative times negative = negative.

OpenStudy (anonymous):

So, it would be 3 times sqrt 2?

zepdrix (zepdrix):

Gotta use brackets in calculator: \(\Large\rm -3^2\) is not the same as \(\Large\rm (-3)^2\) silly gal! :d

OpenStudy (anonymous):

You're right. Hehe, I should probably check my other problems too. I keep forgetting, silly me.

zepdrix (zepdrix):

3sqrt2? Hmm I came up with something different. Lemme see if I did that correctly.

zepdrix (zepdrix):

\[\Large\rm r=\sqrt{(-3)^2+(3\sqrt3)^2}\]\[\Large\rm r=\sqrt{9+9(3)}\]\[\Large\rm r=\sqrt{35}\]

zepdrix (zepdrix):

Did you remember to square both the 3 and the sqrt3 in the second term? :o

OpenStudy (anonymous):

Oh! Geeze, this is embarrassing. I didn't.

zepdrix (zepdrix):

Err I guess it's sqrt(36) right? My addition was a bit off.

OpenStudy (anonymous):

Yeah, because sqrt 35 gives a totally different answer than the choices given to me. So that eliminates two of the answer choices.

OpenStudy (anonymous):

The next thing we're looking for is theta, correct?

zepdrix (zepdrix):

\[\Large\rm r=6\]Yes, correct.

OpenStudy (anonymous):

So for theta, we have to use.. \[\tan^{-1} \frac{ b }{ a }\]

zepdrix (zepdrix):

Yes, good. I would probably just leave it like this though, and relate it to your unit circle.\[\Large\rm \tan \theta=\frac{b}{a}\] But I mean, yes that's how you explicitly solve for theta.

OpenStudy (anonymous):

so it would look like: \[\tan \theta = 3\sqrt{3}/3 \]

zepdrix (zepdrix):

With a negative in there somewhere, yes.\[\Large\rm \tan \theta=-\sqrt3\]Simplify!! :O

OpenStudy (anonymous):

Okay, so going off of common sense. The only time that tan is ever negative is in the second and fourth quadrant.

zepdrix (zepdrix):

Which angle will we want, the one in the `second quadrant` or the `fourth quadrant` angle? `Hint`: Look back your at expression to see which quadrant we're in! \(\Large\rm -3+3\sqrt3\mathcal i\)

OpenStudy (anonymous):

I'll assume second quadrant.

zepdrix (zepdrix):

Real component is negative, imaginary is positive, So yah that puts us up in the second quadrant.

zepdrix (zepdrix):

Do you remember your tangent special angles? :O Those are a little harder to remember than sine and cosine I think.

OpenStudy (anonymous):

I had to look it up, but 60 degrees which is pi/3 has a tangent of sqrt 3.

OpenStudy (anonymous):

So, it would be 2pi/3 for this problem.

zepdrix (zepdrix):

Mmm ok good, that sounds right!

OpenStudy (anonymous):

Great! Thank you so much!

zepdrix (zepdrix):

\[\Large\rm r=6,\qquad \theta=\frac{2\pi}{3}\]Yay team \c:/